OpenStudy (anonymous):

A cyclist rides his bike at a rate of 27 miles per hour. What is this rate in feet per second? How many feet will the cyclist travel in 15 seconds? In your computations, use the fact that 1 mile is equal to 5280 feet. Do not round your answers.

OpenStudy (anonymous):

type "27 mi/hr to ft/s" in google :D

OpenStudy (anonymous):

Ok! but how do I find the 2nd part ?

OpenStudy (anonymous):

distance = speed * time

OpenStudy (anonymous):

Ok so the distance= 39.6 (ft/s) * 15 (seconds) ?

OpenStudy (anonymous):

Sorry I suck at math lol

OpenStudy (***[isuru]***):

yes... u just have to multiply those values :)

OpenStudy (anonymous):

27 miles=27*1760 yards=27*1760*3 feet 1 hour =1*60 minutes=1*60*60 seconds \[27 m/hr=\frac{ 27*1760*3 }{ 1*60*60}=39.6 feet/s\] distance=speed*time calculate.